OpenStudy (anonymous):

a dartboard consists of 2 concentric circles. the probability of hitting the inner circle is 16%. if the outer circle has a diameter of 20 inches. what is the radius of the inner circle?

jimthompson5910 (jim_thompson5910):

we're given this 

jimthompson5910 (jim_thompson5910):

the first step is to find the area of the outer circle

jimthompson5910 (jim_thompson5910):

use A = pi*r^2

OpenStudy (anonymous):

a=31.4

jimthompson5910 (jim_thompson5910):

no it's not r times 2 it's r squared

jimthompson5910 (jim_thompson5910):

A = pi*r^2 A = 3.14*10^2 A = 314.159265

OpenStudy (anonymous):

right 10*10

jimthompson5910 (jim_thompson5910):

that's the area of the larger circle

jimthompson5910 (jim_thompson5910):

sry, if you're just using 3.14 for pi, then it's 314

jimthompson5910 (jim_thompson5910):

now if the probability of hitting the inner circle is 16%, then the are of the smaller circle is 16% of the area of the larger circle

jimthompson5910 (jim_thompson5910):

Area of Smaller Circle = 16% of Area of Larger Circle Area of Smaller Circle = 16% of 314 Area of Smaller Circle = 0.16 * 314 Area of Smaller Circle = ???

OpenStudy (anonymous):

16% of 314=50.24

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now use this to solve for r A = pi*r^2 50.24 = 3.14*r^2 keep going to find r

OpenStudy (anonymous):

got it working on it

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

that's your final answer

OpenStudy (anonymous):

so simple thanks for explaining

jimthompson5910 (jim_thompson5910):

yw